Klints Inc. paid an annual dividend of $1.25 last year. The firm's stock sells for $30.50 per share, and the company is expected to grow at about 4% per year into the foreseeable future. Estimate Klints' cost of retained earnings. Round the answer to two decimal places. Do not round intermediate calculations. ___% _____________________________

Harris Inc.'s preferred stock was issued five years ago to yield 9%. Investors buying those shares on the secondary market today are getting a 14% return. Harris generally pays flotation costs of 13.5% on new securities issues. What is Harris's cost of preferred financing? Round the answer to two decimal places. ____%

explain please
Answer #1

Price = D1 = $30.50 [ 1.25*1.04] D1/(ke-g) = $1.30 4.00% 4.26% 8.26% ke-g = ( 1.30/30.50] (4.0%+4.26% ] ke= cost of retained